OpenStudy (anonymous):

how do i simplify -5i-(-4+6i) and write it in standard form?

OpenStudy (kropot72):

Group like terms first. Can you do that?

OpenStudy (anonymous):

multiply the once in bracket by - it gives you -5i+4-6i collect like terms;we have -5i-6i+4 -11i+4

OpenStudy (anonymous):

actually the final answer is 4-11i but thank you

OpenStudy (kropot72):

So in standard form it is 4-11i

OpenStudy (anonymous):

okay so what about (-2+i)(-3-4i)

OpenStudy (kropot72):

Is the answer 10+5i? If so shall I run thru the solution?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

please explain

OpenStudy (kropot72):

When you multiply the two terms you get: 6-3i+8i+4. Remember that i^2=-1. Can you follow the step above?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank u

OpenStudy (kropot72):

Then you group like terms: (6+4) +(-3i+8i)=10+5i